# An Analysis of Decision Problems for Relational Pattern Languages under Various Constraints

### v1

@@ Appendix 0.A Additional Content for Membership Problem
$h(\alpha_{2,j})=h(z_{j,1}\ f_{j}(X_{j,1})\ z_{j,2}\ f_{j}(X_{j,2})\ z_{i,3}\ f_{j}(X_{j,3})\ z_{j,4})=1^{10}\#1^{10}\#1^{10}=t_{j}.$
By the first equality, we obtain that either $h(u_{i,1})=1$ and $h(v_{i,1})=\#1$ or $h(u_{i,1})=1\#$ and $h(v_{i,1})=1$. Notice by the definition of $r$, in this case, that for all $i\in[m]$ and $j\in[n]$ we have that $h(u_{i,j})\in\{1\}^{*}$ if and only if $h(u_{i,0})=1$. The same holds analogously for all $v_{i,j}$. Now select an assignment for the variables in $\mathcal{X}$ for the $3\mathtt{-SAT}$ formula by setting $X_{i}=true$ (hence $\overline{X_{i}}=false$) iff $h(u_{i,1})=1$ and otherwise $\overline{X_{i}}=true$ (hence $X_{i}=false$) iff $h(v_{i,1})=1$. By the argument from before, this is a valid assignment of the variables. If there were only falsely assigned variables in a single clause, using the previous observations, there existed $j\in[n]$ such that $|h(\alpha_{2,j})|_{\#}\geq 3$, which is not the case. Hence, there is at least one variable assigned to true per clause and by that we obtain that $\varphi$ must be satisfiable. ∎
Proof of Lemma˜2:
Proof
Again, the original proof in [Jiang1994] for the case of equality between variables uses a reduction from $3\mathtt{-SAT}$ to the membership problem. The construction is essentially the same as in [Angluin1980] just with different sizes of each part of the word $w$ to construct. Here, we have for all $i\in[m]$ and $j\in[n]$ that $s_{i}=1$, $t_{j}=1^{7}$, and $w_{j}=1^{2}$. By that, as variables are allowed to be substituted by the empty word, the reduction works in exactly the same way. Consider [Angluin1980, Jiang1994] for the formal proof of this construction. As the reasoning about the construction works exactly in the same way as before, and as the considered relations don’t change in their general behavior if the empty word is allowed, we can use the same arguments as in Lemma 1 and obtain NP-hardness for all these cases here as well. ∎
Proof of Proposition˜3:
Proof
Consider the construction from Proposition 2. As mentioned before, because of the definition of $r_{com^{+}}$, two related variables may never be substituted to the empty word, as $\varepsilon\notin\{y\}^{+}$ for all $y\in\Sigma^{*}$. As only the variables with the base name $z$ are independent, only those may be substituted emptily. Those are only used to fill up remaining #\#’s or $1$’s in the encoded clauses in the case of valid assignments for the $3\mathtt{-SAT}$ formula $\varphi$. Thus, the main argument stays the same and can be applied here in exactly the same way. ∎
Proof of Proposition˜4:
Proof
Again, we reduce from $3\mathtt{-SAT}$. Let $\varphi$ be a $3\mathtt{-SAT}$ formula, let $\mathcal{X}=\{X_{1},...,X_{m}\}$ be the variables in $\varphi$ for some $m\in\mathbb{N}$, let $C=\{C_{1},...,C_{n}\}$ be the set of clauses in $\varphi$ for some $n\in\mathbb{N}$. For each clause $C_{j}\in C$, refer to its three variables by $X_{j,1}$, $X_{j,2}$, and $X_{j,3}$ for $j\in[n]$. W.l.o.g. assume that no variable occurs two times in the same clause (a variable and its reversal, however, may still occur together in one clause). Define $n+1$ functions $f_{j}$ with $j\in[n]_{0}$ such that $f_{j}:\mathcal{X}\rightarrow X$ with $f_{j}(X_{i})=u_{i,j}$ and $f_{j}(\overline{X_{i}})=v_{i,j}$ for $i\in[m]$. Now, let
$w=\#\#\ s_{1}\ \#\#\ ...\ \#\#\ s_{m}\ \#\#\ t_{1}\ \#\#\ ...\ \#\#\ t_{n}\ \#\#$
with
$s_{i}$ $=1\#1,\text{ and}$
$t_{j}$ $=1$
for $i\in[m]$ and $j\in[n]$ and construct a relational pattern $(\alpha,r_{\alpha})$ by
$\alpha=\#\#\alpha_{1,1}\ \#\#\ ...\ \#\#\ \alpha_{1,m}\ \#\#\ \alpha_{2,1}\ \#\#\ ...\ \#\#\ \alpha_{2,n}\ \#\#$
with
$\alpha_{1,i}$ $=f_{0}(X_{i})f_{0}(\overline{X_{i}})\text{, and}$
$\alpha_{2,j}$ $=f_{j}(X_{j,1})f_{j}(X_{j,2})f_{j}(X_{j,3})$
and the set of related variables $r_{\alpha}=\{\ (r,u_{i,j},u_{i,j^{\prime}}),(r,v_{i,j},v_{i,j^{\prime}})\ |\ i\in[m],j,j^{\prime}\in[n+1]\ \}$. Hence, notice that two variables with the same base name $u_{i,j}$ and $u_{i^{\prime},j^{\prime}}$ (resp. $v_{i,j}$ and $v_{i^{\prime},j^{\prime}}$) are related to each other if and only if $i=i^{\prime}$.
First, assume there exists a satisfying assignment for $\varphi$. Let $h\in H$ such that for all $i\in[m]$ and $j\in[n]_{0}$ we have $h(u_{i,0})=1$ and $h(v_{i,0})=\#1$ iff $X_{i}$ is true (hence $\overline{X_{i}}$ is false) and otherwise $h(u_{i,0})=1\#$ and $h(v_{i,0})=1$ iff $X_{i}$ is false (hence $\overline{X_{i}}$ is true). As the assignment for $\varphi$ is true, we have for each clause that there is at least one variable $u_{i,j}$ (or $v_{i,j}$) such that $(r_{com^{*}},u_{i,j},u_{i,0})\in r_{\alpha}$ (or $(r_{com^{*}},v_{i,j},v_{i,0})$) and $h(u_{i,0})=1$ (or $h(v_{i,0})=1$). Hence we may set $h(u_{i,j})=1$ (or $v_{i,j}=1$) and for all other variables $z$ in that clause set $h(z)=\varepsilon$. As the empty word commutes with any other word by the definition of $r_{com^{*}}$, this substitution is $r_{\alpha}$-valid. As we can do that for each $j\in[n]$, we get can find a $r_{\alpha}$-valid substitution $h\in H$ such that $h(\alpha)=w$, concluding this direction.
For the other direction assume $w\in L_{E}(\alpha,r_{\alpha})$. Then there exists some $r_{\alpha}$-valid substitution $h\in H$ such that $h(\alpha)=w$. For each $j\in[n]$, we know that $h(f_{j}(X_{j,1})f_{j}(X_{j,2})f_{j}(X_{j,3}))=1$. Hence, w.l.o.g. there exists one variable $u_{i,j}$ for $i\in[m]$, such that $h(u_{i,j})=1$ and $(r_{com^{*}},u_{i,1},u_{i,j})\in r_{\alpha}$. By the definition of $r_{com^{*}}$ and $w$ we know that $h(u_{i,0})\in\{\varepsilon,1\}$ for $h$ to be $r_{\alpha}$-valid. In that case we know that $h(v_{i,0})\in\{1\#1,\#1\}$ and by that we have $h(v_{i,j})\neq 1$ for any $j\in[n]$ by the definition of $r_{\alpha}$. So, we can set an assignment of $X$ for $\varphi$ by setting $X_{i}=true$ iff $h(u_{i,1})\in\{\varepsilon,1\}$ and setting $\overline{X_{1}}=true$ iff $h(v_{i,1})\in\{\varepsilon,1\}$. As for each clause $j$ we find either a variable $u_{i,j}$ or $v_{i,j}$ substituted to $1$, we know that the assignment for $X$ is satisfying for $\varphi$. That concludes the reduction. ∎
0.A.1 The Membership Problem under the One-Sided Star and One-Sided Subsequence Relations
In this subsection, we briefly discuss the special cases mentioned in the end of [Angluin1980], which considers one-sided relations of potential interest. These cases cannot be expressed directly using the framework of this paper.
Consider, for example, the pattern $\alpha=\mathtt{a}x\mathtt{b}x^{*}\mathtt{c}$ that assumes that $h(x^{*})\in\{h(x)\}^{*}$ needs to hold for some valid substitution $h\in H$. This can be mapped using the relation $r_{*}$, but the way Angluin described that case, we should not be able to have the $r_{*}$ hold in both directions between $x$ and $x^{*}$, which, however, is allowed in relational pattern languages over $r_{*}$. Hence, these special cases of one-sided relations are considered for the subsequence and the star relation at the end of this section.
For simplicity, only in this subsection, we denote patterns similar to the notation used in [Angluin1980], i.e., for the one-sided star we write, for example, $\alpha=\#x\#x^{*}\#$ instead of introducing some variables $x_{1}$ and $x_{2}$ separately and relating them externally by some relation $r$. In particular, in this subsection, given $x$ and $x^{*}$, a substitution $h$ is valid if and only if $h(x^{*})\in\{x\}^{*}$, and, given $x$ and $x^{s}$, a substitution $h$ is valid if and only if $h(x^{s})\in\operatorname{SubSeq}(x)$. For the erasing case, we obtain that the same reduction can be used for both cases, one-sided star and one-sided subsequence.
Proposition 6
Given $w\in\Sigma^{*}$ and some pattern $\alpha$ where each variable $x$ occurs at most once and, if it does, $x^{*}$ (resp. $x^{s}$) may occur multiple times, it is NP-complete to decide whether $w\in L_{E}(\alpha)$.
Proof
NP-containment follows from the fact that we can guess a valid substitution $h$ that results in $|h(\alpha)|=|w|$ and check whether $h(\alpha)=w$ holds, all in polynomial time. To establish NP-hardness, we reduce from $3\mathtt{-SAT}$ as before and use the same formal framework as used in the reductions before. Assuming the number of variables in the formula is $m\in\mathbb{N}$ and assuming the number of clauses of the formula is $n\in\mathbb{N}$, the word $w\in\Sigma^{*}$ will be set to
$w=\#\#\underbrace{\ 1\ \#\ ...\ \#\ 1\ }_{m\text{ blocks}}\#\#\underbrace{\ 1\ \#\ ...\ \#\ 1\ }_{n\text{ blocks}}\#\#.$
The corresponding pattern $\alpha$ is given by
$\alpha=\#\#\ \alpha_{1,1}\ \#\ ...\ \#\ \alpha_{1,m}\ \#\#\ \alpha_{2,1}\ \#\ ...\ \#\ \alpha_{2,n}\ \#\#$
where $\alpha_{1,i}=u_{i}v_{i}$, for $i\in[m]$ and $u_{i},v_{i}$ being new pattern variables, and $\alpha_{2,i}=f(X_{i,1})f(X_{i,2})f(X_{i,3})$, where $f(X_{i,j})=u_{k}^{*}$ if $X_{i,j}=X_{k}$ (resp. $f(X_{i,j})=v_{k}^{*}$ if $X_{i,j}=\overline{X_{k}}$) in the one-sided star framework, and $f(X_{i,j})=u_{k}^{s}$ if $X_{i,j}=X_{k}$ (resp. $f(X_{i,j})=v_{k}^{s}$ if $X_{i,j}=\overline{X_{k}}$) in the one-sided subsequence framework.
The correctness of the reduction comes from the fact that, for $w\in L_{E}(\alpha)$ to hold, exactly one variable in each $\alpha_{1,i}$, for $i\in[m]$, must be set to $1$, either $h(u_{i})=1$ or $h(v_{i})=1$. A variable $u_{i}^{*}$ or $v_{i}^{*}$ (resp. $u_{i}^{s}$ or $v_{i}^{s}$), for $i\in[m]$, that occurs in some $\alpha_{2,j}$, for $j\in[m]$, can only be substituted by $1$ if the corresponding variable in $u_{i}$ or $v_{i}$ in $\alpha_{1,i}$ is substituted by $1$. Bringing these two facts together immediately results in that the $3\mathtt{-SAT}$ formula must have an accepting configuration if and only if $w\in L_{E}(\alpha)$, as we can only substitute all $\alpha_{2,i}$, for $i\in[n]$, by $1$ if and only if there exists some configuration of the variables where either $X_{k}$ or $\overline{X_{k}}$ is set to true and each clause contains at least one of these variables that are set to true. ∎
For the non-erasing cases, we can again use similar reductions as for all the other problems before, but the configuration is slightly different for one-sided star and one-sided subsequence, respectively.
Proposition 7
Given $w\in\Sigma^{*}$ and some pattern $\alpha$ where each variable $x$ occurs at most once and, if it does, $x^{*}$ may occur multiple times, it is NP-complete to decide whether $w\in L_{NE}(\alpha)$.
Proof
NP-containment follows from the fact that we can guess a valid substitution $h$ that results in $|h(\alpha)|=|w|$ and check whether $h(\alpha)=w$ holds, all in polynomial time. To establish NP-hardness, we reduce from $3\mathtt{-SAT}$ as before, again, assuming that a given formula $\phi$ has $n$ clauses and uses $m$ distinct variables. We set $w\in\Sigma^{*}$ to
$w=\#\#\underbrace{\ 1^{3}\ \#\ ...\ \#\ 1^{3}\ }_{m\text{ blocks}}\#\#\underbrace{\ 1^{6}\ \#\ ...\ \#\ 1^{7}\ }_{n\text{ blocks}}\#\#$
The pattern $\alpha$ is basically a prefix of the construction used in Lemma 1, i.e.,
$\alpha=\#\#\ p_{1}\ \#\ ...\ \#\ p_{m}\ \#\#\ q_{1}\ \#\ ...\ \#\ q_{n}\ \#\#,$
where $p_{1}$ to $p_{m}$ and $q_{1}$ to $q_{m}$ are constructed as in the proof of Lemma 1, just with the only difference that every occurrence of $u_{i}$ or $v_{i}$, for $i\in[m]$, in some $q_{1}$ to $q_{n}$, is replaced by $u_{i}^{*}$ and $v_{i}^{*}$ respectively. Then, each $u_{i}$ and $v_{i}$ without a start occurs only once, that is in $p_{i}$. All other occurrences are with a star. The variables $z_{i}$, for $i\in[n]$, are unchanged in $q_{i}$ to $q_{n}$.
To show correctness of this reduction, notice the fact that each $p_{i}$ is still to be substituted to $1^{3}$, resulting in one of $u_{i}$ or $v_{i}$ being substituted by $1$ while the other is substituted by $11$. Assume that a substitution to $1$ represents a true assignment while a substitution by $11$ represents a false assignment. For each clause $i\in[n]$, we know that $q_{i}=x_{i,1}^{*}x_{i,2}^{*}x_{i,3}^{*}z_{i}$, assuming each variable $x_{i,j}$ is either some $u_{k}$ or $v_{k}$, for $k\in[m]$. We know that $z_{i}$ must be substituted by a word of at least length $1$. Hence, for $h(q_{i})=1^{6}$ to be the case, we then need $|h(x_{i,1}^{*}x_{i,2}^{*}x_{i,3}^{*})|\leq 5$. If $h(x_{i,j})=11$, for all $j\in[3]$, then we know that $|h(x_{i,1}^{*}x_{i,2}^{*}x_{i,3}^{*})|\geq 6$, as $h(x_{i,j})=\varepsilon$ is not valid in the non-erasing case. So, for at least one variable $x_{i,j}$ we need $h(x_{i,j})=1$ for $h(q_{i})=1^{6}$ to be possible. As this holds for all clauses, we can only find a substitution $h$ for which $h(\alpha)=w$ if and only if the given $3\mathtt{-SAT}$ is satisfiable. ∎
As we only do slight changed to the constructed word and pattern, we give the proof of the following result in a very short form and assume the same formal framework as in Proposition 7.
Proposition 8
Given $w\in\Sigma^{*}$ and some pattern $\alpha$ where each variable $x$ occurs at most once and, if it does, $x^{s}$ may occur multiple times, it is NP-complete to decide whether $w\in L_{NE}(\alpha)$.
Proof
NP-containment follows as before, e.g. as in Proposition 7. For NP-hardness, we reduce from $3\mathtt{-SAT}$ as before. The word $w\in\Sigma^{*}$ is given by
$w=\#\#\underbrace{\ 1^{3}\ \#\ ...\ \#\ 1^{3}\ }_{m\text{ blocks}}\#\#\underbrace{\ 1^{4}\ \#\ ...\ \#\ 1^{4}\ }_{n\text{ blocks}}\#\#.$
The pattern $\alpha$ is given by
$\alpha=\#\#\ p_{1}\ \#\ ...\ \#\ p_{m}\ \#\#\ q_{1}\ \#\ ...\ \#\ q_{n}\ \#\#,$
where $p_{i}=u_{i}v_{i}$, for $i\in[m]$ and new variables $u_{i}$ and $v_{i}$, and we have $q_{i}=f(X_{i,1})f(X_{i,2})f(X_{i,3})$ for $f$ being defined as before as well, just using $u_{j}^{s}$ and $v_{j}^{s}$ instead of $u_{j}$ and $v_{j}$, for $j\in[m]$ and $i\in[n]$. Now notice, again, that for each $p_{i}$, for $i\in[m]$, one of $u_{i}$ and $v_{i}$ must be substituted by $1$ while the other one is substituted by $11$. Here, we consider a substitution by $11$ as a true assignment of that variable configuration. For all other variable substitutions we know that they must be a subsequence of either some $u_{i}$ or $v_{i}$. As we are in the non-erasing case, for a substitution $h$ to be valid, that results in w.l.o.g. $h(u_{i}^{s})=1$ if $h(u_{i})=1$ and $h(u_{i}^{s})\in\{1,11\}$ if $h(u_{i})=11$, for $i\in[m]$ (respectively for $v_{i}$). So, we can only obtain $h(q_{i})=1^{4}$, for all $i\in[n]$, if and only if there is at least one variable $u_{i}^{s}$ (or $v_{i}^{s}$) in $q_{i}$, where $h(u_{i})=11$ (or $h(v_{i})=11$). By the same arguments as in all results before, we can only find a substitution $h\in H$ for which $h(\alpha)=w$ if and only if there exists a satisfying assignment of the variable for the given $3\mathtt{-SAT}$ formula, concluding this reduction. ∎
This concludes all results necessary for Theorem 3.2 to hold.
Appendix 0.B Additional Content for Inclusion Problem
0.B.1 Definition of Nondeterminstic 2-Counter Automata without Input
A nondeterministic 2-counter automaton without input (see e.g. [Iberra1978]) is a 4-tuple $A=(Q,\delta,q_{0},F)$ which consists of a set of states $Q$, a transition function $\delta:Q\times\{0,1\}^{2}\rightarrow\mathcal{P}(Q\times\{\-1,0,+1\}^{2})$, an initial state $q_{0}\in Q$, and a set of accepting states $F\subseteq Q$. A configuration of $A$ is defined as a triple $(q,m_{1},m_{2})\in Q\times\mathbb{N}\times\mathbb{N}$ in which $q$ indicates the current state and $m_{1}$ and $m_{2}$ indicate the contents of the first and second counter. We define the relation $\vdash_{A}$ on $Q\times\mathbb{N}\times\mathbb{N}$ by $\delta$ as follows. For two configurations $(p,m_{1},m_{2})$ and $(q,n_{1},n_{2})$ we say that $(p,m_{1},m_{2})\vdash_{A}(q,n_{1},n_{2})$ if and only if there exist $c_{1},c_{2}\in\{0,1\}$ and $r_{1},r_{2}\in\{-1,0,+1\}$ such that
  1. 1.
    if $m_{i}=0$ then $c_{i}=0$, otherwise if $m_{i}>0$, then $c_{i}=1$, for $i\in\{1,2\}$,
  2. 2.
    $n_{i}=m_{i}+r_{i}$ for $i\in\{1,2\}$,
  3. 3.
    $(q,r_{1},r_{2})\in\delta(p,c_{1},c_{2})$, and
  4. 4.
    we assume if $c_{i}=0$ then $r_{i}\neq-1$ for $i\in\{1,2\}$.
and let $\mathtt{InvalC}(A)=\{0,\#\}^{*}\setminus\mathtt{ValC}(A)$. The emptiness problem for deterministic 2-counter-automata is undecidable (cf. e.g. [Iberra1978, Minsky1961]), thus it is also undecidable whether a nondeterministic 2-counter automaton without input has an accepting computation [Freydenberger2010, Jiang1995]. That the emptiness problem for universal Turing machines is undecidable is a known fact.
0.B.2 Definition of the Universal Turing Machine $U$
Here, we define the universal Turing machine $U$ used in the referenced proof from [BREMER201215] and referred to in the proofs for the undecidability of the inclusion problem for NE relational pattern languages under $r_{ab}$ and rev in Proposition 5 and Corollary 2. Let $U=(Q,\Gamma,\delta)$ be the universal Turing machine $U_{15,2}$ wih 2 symbols and 15 states as described by Neary and Woods [Neary2009]. This machine hat the states $Q=\{q_{1},...,q_{15}\}$ and the tape alphabet $\Gamma=\{0,1\}$. The transition function $\delta:\Gamma\times Q\rightarrow(\Gamma\times\{L,R\}\times Q)\cup\mathtt{HALT}$ is given in Table 4.
We follow with the definition of encodings of computations of $U$ as depicted in [BREMER201215]. The following conventions are needed to discuss configurations of $U$. The tape content of any configuration of $U$ is characterized by two infinite sequences $t_{L}=(t_{L,n})_{n\geq 0}$ and $t_{R}=(r_{R,n})_{n\geq 0}$ over $\Gamma$. The sequence $t_{L}$ describes the left side of the tape, the sequence starting at the head position of $U$ (including) and extending to the left. Analogously, $t_{R}$ describes the right side of the tape, the sequence starting directly after the head position and extending to the right. A configuration $C=(q_{i},t_{L},t_{R})$ of $U$ is a triple consisting of a state $q_{i}$, a left side of the tape $t_{L}$ and a right side of the tape $t_{R}$.

### v2

@@ 0.A.5 Proof of Lemma˜3
$\mathbf{occ}_{\alpha,i}(j)\geq\sum_{C_{\alpha,i^{\prime}}\in S^{\prime}}\mathbf{occ}_{\alpha,i^{\prime}}(j).$
Then all words generated in each block individually by $C_{\alpha,i}$ can also be generated by the variables in $S^{\prime}$ and we may remove all variables of that equivalence class without changing the language. We may consider each block individually since $c$ does not occur as a terminal, we may assign the letter $c$ to all variable occurrences. By iteratively removing dominated classes we obtain a minimal pattern of $\alpha$. For a removed class $i$, we set $\mathbf{occ}_{\alpha,i}(j)=0$ for all $j\in\{0,\dots,m\}$. We do the same reduction step for $\beta$. Now, the languages are equal if and only if the minimal patterns are identical, i.e., for all $i\in\{1,\dots,k_{\alpha}\}$ there exists $i^{\prime}\in\{1,\dots,k_{\beta}\}$ such that $\mathbf{occ}_{\alpha,i}=\mathbf{occ}_{\beta,i^{\prime}}$.
We now argue that this condition is correct. Suppose the minimal patterns differ. Then there exists a class $C$ for one pattern that has no matching class in the other pattern but is also not dominated. Because $c$ is unused as a terminal, construct a word that assigns all variables of $C$ the letter $c$ and sets all other variables to $\varepsilon$. Due to the distinct number of occurrences, this gives in at least one block a word length that can not be matched by the other pattern. Thus, the languages differ.
The block decomposition and counting the occurrences both require linear time. Testing the dominance between classes takes at most $O(k_{\alpha}^{2}m)$, and respectively $O(k_{\beta}^{2}m)$ comparisons. Comparing the minimal patterns is also possible in linear time Hence, the complete procedure runs in polynomial time. ∎
0.A.6 Proof of Proposition˜1
Proof.
For $r\in\{r_{*},r_{ssq}\}$, reduce from equivalence of erasing pattern languages over $r_{=}$ as follows. Given $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ with $R=\{r_{=}\}$, construct $(\alpha,r^{\prime}_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ with $R=\{r\}$ by replacing every $(r_{=},x_{i},x_{j})\in r_{\alpha}$ with both $(r,x_{i},x_{j})$ and $(r,x_{j},x_{i})$. Since symmetric $r_{*}$-constraints and symmetric $r_{ssq}$-constraints force equality of substituted words, this preserves exactly the valid substitutions. Hence the equivalence problem for erasing pattern languages reduces directly to equivalence for relational patterns over $r_{*}$ and $r_{ssq}$.
For $r=r_{rev}$, let $(\alpha,r_{\alpha})\in\mathtt{RelPat}_{\Sigma,R}$ with $R=\{r_{=}\}$. Define
$\overset{\leftrightarrow}{\alpha}=\alpha\overset{\leftarrow}{\alpha},$
where $\overset{\leftarrow}{\alpha}$ is the reversal of $\alpha$ in which each variable $x_{i}$ is replaced by a fresh variable $x_{i^{\prime}}$. Let $r_{\overset{\leftrightarrow}{\alpha}}$ contain $(r_{rev},x_{i},x_{i^{\prime}})$ for every $x_{i}\in\operatorname{var}(\alpha)$ and, for every $(r_{=},x_{i},x_{j})\in r_{\alpha}$, also $(r_{rev},x_{i},x_{j^{\prime}})$ and $(r_{rev},x_{j},x_{i^{\prime}})$. Then every valid substitution satisfies $h(x_{i^{\prime}})=h(x_{i})^{R}$, and the additional constraints imply $h(x_{i})=h(x_{j})$ iff $(r_{=},x_{i},x_{j})\in r_{\alpha}$. Consequently, $L_{E}(\overset{\leftrightarrow}{\alpha},r_{\overset{\leftrightarrow}{\alpha}})=\{ww^{R}\mid w\in L_{E}(\alpha,r_{\alpha})\}.$ Thus, for $(\alpha,r_{\alpha}),(\beta,r_{\beta})\in\mathtt{RelPat}_{\Sigma,R}$ over $r_{=}$, $L_{E}(\alpha,r_{\alpha})=L_{E}(\beta,r_{\beta})$ if and only if $L_{E}(\overset{\leftrightarrow}{\alpha},r_{\overset{\leftrightarrow}{\alpha}})=L_{E}(\overset{\leftrightarrow}{\beta},r_{\overset{\leftrightarrow}{\beta}}).$ Hence equivalence of erasing pattern languages reduces to the equivalence problem for relational patterns over each $r\in\{r_{ssq},r_{*},r_{rev}\}$. ∎
Appendix 0.B Additional Content for the Inclusion Problem
0.B.1 Proof Sketch of Theorem˜4.1
0.B.1.1 (A Specific Universal Turing Machine):
Here, we define the universal Turing machine $U$ used in the referenced proof from [5] and referred to in the proofs for the undecidability of the inclusion problem for NE relational pattern languages under $r_{ab}$ and rev in this section. Let $U=(Q,\Gamma,\delta)$ be the universal Turing machine $U_{15,2}$ wih 2 symbols and 15 states as described by Neary and Woods [31]. This machine hat the states $Q=\{q_{1},...,q_{15}\}$ and the tape alphabet $\Gamma=\{0,1\}$. The transition function $\delta:\Gamma\times Q\rightarrow(\Gamma\times\{L,R\}\times Q)\cup\mathtt{HALT}$ is given in Table 2.
We follow with the definition of encodings of computations of $U$ as depicted in [5]. The following conventions are needed to discuss configurations of $U$. The tape content of any configuration of $U$ is characterized by two infinite sequences $t_{L}=(t_{L,n})_{n\geq 0}$ and $t_{R}=(r_{R,n})_{n\geq 0}$ over $\Gamma$. The sequence $t_{L}$ describes the left side of the tape, the sequence starting at the head position of $U$ (including) and extending to the left. Analogously, $t_{R}$ describes the right side of the tape, the sequence starting directly after the head position and extending to the right. A configuration $C=(q_{i},t_{L},t_{R})$ of $U$ is a triple consisting of a state $q_{i}$, a left side of the tape $t_{L}$ and a right side of the tape $t_{R}$.